OpenStudy (anonymous):

question 1 find the midpoint of the segment with endpoint (5,2) and (-1,8)

OpenStudy (anonymous):

(5+-1)/2 is the x-coordinate of the midpoint and (2+8)/2 is the y-coordinate of the midpoint it's almost like averaging the points to fin d the midpoint ;)

OpenStudy (anonymous):

but how i dont realy get it

OpenStudy (aravindg):

miidpoint formula=((x1+x2)/2,(y1+y2)/2) use that !

OpenStudy (anonymous):

so thats the answere

OpenStudy (anonymous):

(2, 5) is the midpoint

OpenStudy (accessdenied):

 If you want the number between two other numbers, you have to essentially add the numbers together and divide the sum by 2. Like, if you wanted the number between 3 and 5 (which is obviously 4), we could take 5 + 3 = 8 and then divide by 2 8/2 = 4. (x1 + x2)/2 = the midpoint x-value (y1 + y2)/2 similarly for the y-value. If that makes some sense of the formula.